1. An auction house has two appraisers of pecious jewelry. The probability that the older of the two will not be available is 0.33, the probability that the other one wil not be available is 0.27, and the probability that both of them will not be available is 0.19. What is the probability that either or both of them will not be available?
1
Expert's answer
2021-05-17T18:56:59-0400

The probability that the older of the two will not be available is

P(A) = 0.33

The probability that the other one will not be available is

P(B) = 0.27

The probability that both of them will not be available is

"P(A \cap B) = 0.19"

The probability that either or both of them will not be available

"P(A \cup B) = P(A) + P(B) -P(A \cap B) \\

= 0.33 + 0.27 -0.19 \\

= 0.41"

So, the probability that either or both of them will not be available is 0.41.